# CANDLE: Iterative Conceptualization and Instantiation Distillation from Large Language Models for Commonsense Reasoning

### v1

@@ 5.2 Downstream Applications
In this section, we explore the downstream applications of CANDLE. By applying CANDLE to ATOMIC, the distilled conceptualizations and instantiations form a large-scale expansion of the original CSKB, which contains high-quality abstract and concrete commonsense knowledge. Leveraging both types of knowledge as supplementary training data, we enhance various commonsense reasoning models for downstream tasks. Specifically, we utilize distilled conceptualizations in two CSKB conceptualization tasks Wang et al. (2023b), while instantiations are used in generative commonsense inference (COMET; Bosselut et al., 2019) and zero-shot commonsense QA tasks Ma et al. (2021). Due to space constraints, for all tasks, please refer to: Appendix B for dataset descriptions and Appendix C for implementation details.
@@ 5.2.1 CSKB Conceptualization
The CSKB conceptualization task evaluates a model’s ability to conceptualize a CSKB through two binary classification subtasks, which are crucial for performing CSKB conceptualization inference upon concept taxonomies He et al. (2022a); Wang et al. (2023b). The first subtask, event conceptualization, aims to determine whether can be correctly conceptualized using , where is derived by replacing an instance with its linked concept . The second subtask, triple conceptualization, aims to assess the plausibility of a conceptualized triple that represents abstract commonsense knowledge. Accuracy is used as the evaluation metric. Following Wang et al. (2023b), we use the AbstractATOMIC dataset provided by He et al. (2022a) as the evaluation benchmark, which extends ATOMIC Sap et al. (2019a) by including human-annotated conceptualizations and abstract commonsense triples. The annotated data is partitioned by following ATOMIC’s original split of head events.
@@
We compare our distilled models against the following baselines. First, we report the performance of language models when fine-tuned on the training set of AbstractATOMIC, including RoBERTa-Large Liu et al. (2019), DeBERTa-V3-Large He et al. (2023), GPT-2 Radford et al. (2019), LLAMA2 Touvron et al. (2023), Mistral Jiang et al. (2023), and VERA Liu et al. (2023). For fine-tuning LLMs, we adopt LoRA Hu et al. (2022) due to computational power limitations. Additionally, we incorporate semi-supervised learning methods that leverage both human-annotated and unlabeled data in AbstractATOMIC, including PsuedoReasoner Fang et al. (2022) and CAT Wang et al. (2023b). Finally, we include the performance of prompting LLMs, including LLAMA2, Mistral, and ChatGPT. In addition to direct zero-shot prompting, we also experiment with augmenting their performance with five-shot exemplars Wei et al. (2023) and chain-of-thought reasoning Wei et al. (2022).
@@
The results are reported in Table 2. Among all baselines, CAT trained with DeBERTa-v3-large demonstrates superior performance for both tasks compared to others. Within the baselines that use LLMs, LLAMA and Mistral yield the best results after fine-tuning. It’s worth noting that prompting ChatGPT with additional exemplars and chain-of-thought also produces commendable performances while prompting LLAMA and Mistral fail in most scenarios. However, pre-training language models on CANDLE distilled conceptualizations leads to consistently improved results for both tasks. This is evident when comparing the performance of each model with its respective best-performing baseline using the same backbone. For instance, fine-tuned Mistral shows a significant improvement of 1.54% and 2.19% on two tasks compared to fine-tuning on AbstractATOMIC, respectively. Remarkably, the distilled DeBERTa-v3-large surpasses all baselines and achieves state-of-the-art performances. This can be attributed to conceptualizations distilled from CANDLE, which provide the model with a broader understanding of various events and their conceptualizations, thus enhancing its discriminatory capabilities.
5.2.2 Generative Commonsense Inference
Task Setup.
The task of generative commonsense inference modeling (COMET; Bosselut et al., 2019) asks the model to generate commonsense tails based on given head and relation inputs. It serves as a fundamental component for numerous high-level applications that necessitate commonsense reasoning, such as zero-shot commonsense question answering with self-talk Shwartz et al. (2020) and dynamic graph construction Bosselut et al. (2021), narrative reasoning Peng et al. (2022b), and dialogue generation Tu et al. (2022). Improving COMET can potentially benefit other domains that require commonsense understanding. Following Hwang et al. (2021), we use the full test set of ATOMIC as our evaluation benchmark, which comprises 34,689 triples across 23 distinct commonsense relations. To automatically evaluate our model, we utilize a range of metrics, including BLEU Papineni et al. (2002), ROUGE-L Lin (2004), METEOR Lavie and Agarwal (2007), CIDEr Vedantam et al. (2015), and BERTScore Zhang et al. (2020). We also recruit four expert annotators to conduct an expert evaluation of the models’ outputs. They are asked to annotate the plausibility of 200 randomly selected commonsense triple generations under each setting, and the resulting plausibility rates are reported.
Baselines.
For baselines, we use GPT2 and LLAMA2-7B as the backbone models. We train them separately on the training sets of ATOMIC, ATOMIC, ATOMIC10X West et al. (2022), and AbstractATOMIC. These models are then fine-tuned on the training split of ATOMIC and evaluated on the test set. We also include their zero-shot prompting performances. LLAMA2 is additionally evaluated for its prompting performance with five-shot exemplars. We also incorporate ChatGPT as another powerful commonsense generation baseline and provide its performance in zero-shot, five-shot, and chain-of-thought prompting scenarios.
Results and Analysis.
The results are presented in Table 3. Among the baselines, models that are pre-trained on ATOMIC-10X demonstrate the highest expert acceptance rate, surpassing even those pre-trained on AbstractATOMIC. One possible reason is that ATOMIC-10X encompasses a broader range of commonsense relations that are more consistent with the testing set of ATOMIC. However, models that are pre-trained on CANDLE instantiations achieve the highest scores compared to baselines with the same backbone model. For example, the CANDLE distilled LLAMA-7B model exhibits a 1.01% increase in BERTScore and a 3.00% increase in expert-plausibility compared to the best baseline. Even when compared to ChatGPT, CANDLE distilled models outperform it in all automatic metrics while maintaining a high plausibility rate of around 80%. We also observe that CANDLE distilled models consistently outperform those pre-trained on ATOMIC-10X and AbstractATOMIC across all metrics. This highlights the superior benefits of using CANDLE distilled instantiations for training COMET, compared to traditional symbolic knowledge distillation methods or augmenting the original CSKB with abstract commonsense knowledge.
@@ 5.2.2 Generative Commonsense Inference
| DeBERTa-v3-L (CANDLE Distilled) | CANDLE | 81.2 | 69.9 | 80.3 | 65.9 | 78.3 | 74.9 |
@@
| VERA-T5-xxl (CANDLE Distilled) | CANDLE | 73.8 | 64.7 | 77.6 | 59.4 | 71.3 | 69.4 |
@@ 5.2.3 Zero-shot Commonsense QA
The task of zero-shot commonsense QA involves selecting the most plausible option for commonsense questions without any supervision signals from benchmark data. Several methods have been proposed to tackle this task, including those by Shwartz et al. (2020); Bosselut et al. (2021); Kim et al. (2022); Shi et al. (2023). The most effective pipeline, as suggested by Ma et al. (2021), injects commonsense knowledge into language models via fine-tuning on QA pairs synthesized from knowledge in CSKBs. During the fine-tuning process, the head and relation of a triple are transformed into a question using natural language prompts, with the tail serving as the correct answer option. Distractors or negative examples are generated by randomly sampling tails from triples that do not share common keywords with the head. This fine-tuning procedure enhances the model’s knowledge not only for QA benchmarks constructed from CSKBs but also improves its ability to answer unseen commonsense questions in a more generalized manner.
We follow a similar approach to obtain our distilled models by first synthesizing QA pairs from various CSKBS. In addition to directly synthesizing from knowledge triples in ATOMIC, we augment ATOMIC by sampling triples from ATOMIC-10X, AbstractATOMIC, and CANDLE instantiations. The number of sampled triples is the same as in the original ATOMIC dataset. We then synthesize them into QA pairs, following the pipeline introduced by Ma et al. (2021). The resulting QA pairs are used to train different baseline models and CANDLE distilled models. For our distilled models, we utilize QA pairs sourced from CANDLE-instantiation augmented ATOMIC to train a DeBERTa-v3-large model using the marginal ranking loss Ma et al. (2021) and a T5-xxl model Raffel et al. (2020) following the training regime of VERA Liu et al. (2023). Following Ma et al. (2021), we evaluate the performance of the models on the validation split of five commonsense QA datasets: Abductive NLI (aNLI; Bhagavatula et al., 2020), CommonsenseQA (CSQA; Talmor et al., 2019), PhysicalIQA (PIQA; Bisk et al., 2020), SocialIQA (SIQA; Sap et al., 2019b), and WinoGrande (WG; Sakaguchi et al., 2021). Accuracy is used as the evaluation metric.
@@
We compare our distilled models against various baselines. For pre-trained language models, we report the performances of vanilla RoBERTa-Large, DeBERTa-v3-Large, Self-talk Shwartz et al. (2020), COMET-DynaGen Bosselut et al. (2021), SMLM Banerjee and Baral (2020), MICO Su et al. (2022), MR Ma et al. (2021), STL-Adapter Kim et al. (2022), and the previous state-of-the-art CAR Wang et al. (2023a). For MR Ma et al. (2021) and CAR Wang et al. (2023a), DeBERTa-v3-Large is used as the backbone and their performances on ATOMIC-10X and AbstractATOMIC are also reported. For LLMs, we first report the performances of prompting GPT3.5 Brown et al. (2020), ChatGPT, GPT4 OpenAI (2023), LLAMA2, and Mistral in a zero-shot manner. For ChatGPT, advanced zero-shot prompting techniques such as chain-of-thought Wei et al. (2022) and self-consistency chain-of-thought Wang et al. (2023c) are used to assist the model in reasoning. We also train several VERA-T5-xxl baselines on different sets of QA pairs as LLM baselines.
@@
Table 4 presents the results, demonstrating that CANDLE distilled models generalize better than the baselines across several commonsense QA benchmarks. For instance, VERA demonstrates an average improvement of 1.4% compared to the best baseline. This can be attributed to the inclusion of new entities and events in CANDLE instantiations that are absent in other CSKBs, where CANDLE instantiations can aid in answering commonsense questions that require knowledge of these new instances. Furthermore, the distilled DeBERTa-v3-large model outperforms all baselines, including methods utilizing LLMs. By incorporating CANDLE instantiations, the performance of QA models is significantly enhanced, surpassing those utilizing ATOMIC-10X and AbstractATOMIC. This also shows the effectiveness of distilling knowledge with CANDLE in downstream tasks compared to simply applying symbolic knowledge distillation or conceptualization augmentation.
@@ 6 Conclusion
This paper introduces CANDLE, a distillation framework that realizes the chain of conceptualization and instantiation over CSKBs. We demonstrate the efficacy of CANDLE through comprehensive evaluations of the distilled knowledge and its positive impact on downstream tasks. Our research sheds light on leveraging LLMs to enable more robust and generalizable commonsense reasoning.
@@
The major limitation of CANDLE lies in the significant cost of distilling LLMs to obtain substantial knowledge. While the instantiation step of CANDLE utilizes the open foundation model LLAMA2, the conceptualization is still performed by ChatGPT due to the unsatisfactory performance of other LLMs and the superior quality of ChatGPT’s generation. Consequently, a considerable amount of funding is required to distill conceptualizations for CANDLE to function effectively. Future works should address this issue by exploring advanced LLMs or employing stronger open LLMs as the foundation for distilling conceptualizations, potentially utilizing advanced prompting techniques.
Furthermore, it should be noted that CANDLE has only been validated on the ATOMIC dataset. However, CANDLE is not limited to any specific format of commonsense knowledge, allowing it to operate on any CSKB. Future research can address this by extending the evaluation of CANDLE on other CSKBs Nguyen et al. (2023); Speer et al. (2017); Fang et al. (2023), and conducting follow-up experiments to explore their benefits on other downstream tasks or domains.
@@
One of the main ethical concerns in this process is the potential for generating harmful or unethical content from models like ChatGPT and LLAMA2. To address this concern, we recruit four expert annotators, who are graduate and undergraduate students specializing in machine commonsense in natural language processing, to verify the ethics and potential harm of the generated content. A thorough assessment of a random sample has been conducted, and no significant harm has been identified. All training and evaluation datasets used are publicly available and shared under open-access licenses solely for research purposes, aligning with their intended usage. These datasets have been carefully anonymized and desensitized to protect data privacy and confidentiality. The expert annotators involved in this study are fully aware of the annotation protocol and the intended use of their annotations. Their participation in this research is voluntary, and they have agreed to contribute without receiving any compensation. Thus, the authors believe that this paper does not raise any ethical concerns to the best of their knowledge.

### v2

@@ 5.2 Downstream Applications
In this section, we explore the downstream applications of CANDLE. By applying CANDLE to ATOMIC, the distilled conceptualizations and instantiations form a large-scale expansion of the original CSKB, which contains high-quality abstract and concrete commonsense knowledge. Leveraging both types of knowledge as supplementary training data, we enhance various downstream commonsense reasoning models. Specifically, we utilize distilled conceptualizations in the CSKB conceptualization task Wang et al. (2023b), while instantiations are used in generative commonsense inference (COMET; Bosselut et al., 2019) and zero-shot commonsense QA tasks Ma et al. (2021). Due to space constraints, please refer to Appendix B, C, D for task setups, dataset descriptions, and implementation details, respectively.
@@ 5.2.1 CSKB Conceptualization
The CSKB conceptualization task evaluates a model’s ability to conceptualize a CSKB through two binary classification subtasks, which are crucial for performing CSKB conceptualization inference upon concept taxonomies He et al. (2024). The first subtask, event conceptualization, aims to determine whether can be correctly conceptualized using , where is derived by replacing an instance with its linked concept . The second subtask, triple conceptualization, aims to assess the plausibility of a conceptualized triple that represents abstract commonsense knowledge. Accuracy is used as the evaluation metric. Following Wang et al. (2023b), we use the AbstractATOMIC dataset provided by He et al. (2024) as the evaluation benchmark.
@@
We evaluate our distilled models by comparing them against several baselines. These include supervised fine-tuned language models like RoBERTa-Large Liu et al. (2019), DeBERTa-V3-Large He et al. (2023), GPT-2 Radford et al. (2019), LLAMA2 Touvron et al. (2023), Mistral Jiang et al. (2023), and VERA Liu et al. (2023), as well as semi-supervised methods such as PsuedoReasoner Fang et al. (2022) and CAT Wang et al. (2023b). Due to computational power limitations, we utilize LoRA Hu et al. (2022) for fine-tuning LLMs. As additional baselines, we also consider prompting LLMs, including LLAMA2, Mistral, and ChatGPT. We explore both direct zero-shot prompting and alternative methods, such as with five-shot exemplars Wei et al. (2023) and chain-of-thought reasoning Wei et al. (2022).
@@
Table 2 shows the results. CAT trained with DeBERTa-v3-large outperforms all other baselines for both tasks. Among LLMs, LLAMA and Mistral perform well after fine-tuning, but they struggle in prompting scenarios. However, pre-training on CANDLE’s distilled conceptualizations consistently improves results for both tasks. For example, Mistral shows a significant improvement of 1.54% and 2.19% on two tasks compared to directly fine-tuning on AbstractATOMIC. Additionally, the distilled DeBERTa-v3-large surpasses all baseline models and achieves state-of-the-art performance. This can be attributed to the distilled conceptualizations obtained from CANDLE, which grant the model a more comprehensive understanding of conceptualizations and subsequently enhance its discriminatory capabilities.
@@
| DeBERTa-v3-L (CANDLE Distilled) | CANDLE | 81.2↑1.6 | 69.9↑0.6 | 80.3↑1.7 | 65.9↑1.9 | 78.3↑0.1 | 74.9↑1.0 |
@@
| VERA-T5-xxl (CANDLE Distilled) | CANDLE | 73.8↑0.6 | 64.7↑1.7 | 77.6↑0.4 | 59.4↑1.2 | 71.3↑3.2 | 69.4↑1.4 |
5.2.2 Generative Commonsense Inference
Task Setup.
The task of generative commonsense inference modeling (COMET; Bosselut et al., 2019) asks the model to generate commonsense tails based on given head and relation inputs. Following Hwang et al. (2021), we use the full test set of ATOMIC as our evaluation benchmark. We use several automatic metrics for evaluation, including BLEU Papineni et al. (2002), ROUGE-L Lin (2004), METEOR Lavie and Agarwal (2007), CIDEr Vedantam et al. (2015), and BERTScore Zhang et al. (2020). Meanwhile, four expert annotators are recruited to conduct expert evaluations of the generations. They are asked to annotate the plausibility of 200 randomly selected commonsense triple generations under each setting, and the resulting plausibility rates are reported.
Baselines.
For baselines, we separately train GPT2 and LLAMA2-7B on the training sets of ATOMIC, ATOMIC, ATOMIC10X West et al. (2022), and AbstractATOMIC. These models are then fine-tuned on the training split of ATOMIC and evaluated on its test set. We also include their zero-shot prompting performances, with LLAMA2 being evaluated with five-shot exemplars. ChatGPT’s performances under zero-shot, five-shot, and chain-of-thought settings are also reported.
Results and Analysis.
Table 3 shows the results. Among the baselines, models pre-trained on ATOMIC-10X achieve the highest expert acceptance rate, surpassing those trained on AbstractATOMIC. This may be because ATOMIC-10X covers a wider range of commonsense relations consistent with ATOMIC. However, CANDLE distilled models achieve the highest scores compared to baselines with the same backbone model. For example, the CANDLE distilled LLAMA-7B model improves BERTScore by 1.01% and expert-plausibility by 3.00% compared to the best baseline. It also outperforms ChatGPT in all automatic metrics while maintaining a high plausibility rate of around 80%. This emphasizes the advantages of using CANDLE distilled instantiations for COMET training over traditional symbolic knowledge distillation methods or conceptualization augmentation. Interestingly, we also observe that LLAMA2 has a tendency to generate long and contextually rich commonsense knowledge. On the other hand, GPT2, when fine-tuned on ATOMIC-like data, may generate shorter and more concise knowledge, which aligns with the format and length of knowledge in ATOMIC2020, thus achieving better results in automatic evaluations. However, human annotators tend to consider long and contextually rich commonsense statements, generated by LLAMA2, as more plausible.
@@ 5.2.3 Zero-shot Commonsense QA
The task of zero-shot commonsense QA involves selecting the most plausible option for commonsense questions without any supervision signals from benchmark data. We follow the most effective pipeline by Ma et al. (2021), which fine-tune language models on QA pairs synthesized from knowledge in CSKBs. The head and relation of a triple are transformed into a question using natural language prompts, with the tail serving as the correct answer option. Distractors or negative examples are generated by randomly sampling tails from triples that do not share common keywords with the head. In addition to directly synthesizing from knowledge triples in ATOMIC, we augment ATOMIC by sampling triples from ATOMIC-10X, AbstractATOMIC, and CANDLE instantiations. The number of sampled triples is the same as in the original ATOMIC dataset. We then synthesize them into QA pairs to train different baseline models and CANDLE distilled models. For our distilled models, we utilize QA pairs sourced from CANDLE-instantiation augmented ATOMIC to train a DeBERTa-v3-large model using the marginal ranking loss and a T5-xxl model Raffel et al. (2020) following the training regime of VERA. We evaluate the performance of all models on the validation split of Abductive NLI (aNLI; Bhagavatula et al., 2020), CommonsenseQA (CSQA; Talmor et al., 2019), PhysicalIQA (PIQA; Bisk et al., 2020), SocialIQA (SIQA; Sap et al., 2019b), and WinoGrande (WG; Sakaguchi et al., 2021). Accuracy is used as the evaluation metric.
@@
First, we report performances of vanilla RoBERTa-Large, DeBERTa-v3-Large, Self-talk Shwartz et al. (2020), COMET-DynaGen Bosselut et al. (2021), SMLM Banerjee and Baral (2020), MICO Su et al. (2022), MR Ma et al. (2021), STL-Adapter Kim et al. (2022), and the previous state-of-the-art method, CAR Wang et al. (2023a). For MR and CAR, DeBERTa-v3-Large is used as the backbone, and their performances on ATOMIC-10X and AbstractATOMIC are also reported. For LLMs, we report the performances of prompting GPT3.5 Brown et al. (2020), ChatGPT, GPT4 OpenAI (2023), LLAMA2, and Mistral in a zero-shot manner. For ChatGPT, its performances with chain-of-thought Wei et al. (2022) and self-consistency chain-of-thought Wang et al. (2023c) prompting are also reported. We also train several VERA-T5-xxl baselines on different sets of QA pairs as LLM baselines.
@@
Table 4 shows the results, demonstrating that CANDLE distilled models generalize better than the baselines across several commonsense QA benchmarks. For instance, VERA demonstrates an average improvement of 1.4% compared to the best baseline. This can be attributed to the inclusion of new entities and events in CANDLE instantiations that are absent in other CSKBs, where CANDLE instantiations can aid in answering commonsense questions that require knowledge of these new instances. Furthermore, the distilled DeBERTa-v3-large model outperforms all baselines, including methods utilizing LLMs. This also indicates that augmenting with CANDLE distilled instantiations provides a more significant advantage compared to using symbolically distilled or abstract knowledge as training data.
| Critic | Conceptualization | Instantiation |
| 0.0 | 92.3% | 85.5% |
| 0.5 | 94.6% | 91.2% |
| 0.7 | 95.9% | 93.3% |
| 0.9 | 98.3% | 96.7% |
5.3 Analysis
5.3.1 Feasibility of Iterating CANDLE
We first demonstrate the feasibility of iterating the CANDLE framework with more than one round. To do so, we randomly sample 10,000 distilled instantiations from LLAMA2 as the input for the CANDLE framework and execute the framework again, resulting in 200,000 conceptualizations and 200,000 instantiations. Subsequently, we randomly select 300 from each set and annotate them accordingly. The results are shown in Table 5. We observe that iterating the framework produces slightly better results than the first loop. This improvement may be attributed to the fact that the knowledge generated in the initial loop is more easily understood by LLMs compared to the human-annotated data in ATOMIC. Moreover, 58% of the conceptualizations and 44% of the instantiations are novel compared to the first loop. Based on these findings, we believe that our iterative framework is effective, and the iteration process enhances the augmentation of a CSKB through multiple iterations.
| Task | CSKB Concept. | COMET | CSQA |
| Overlap Ratio | 10.1% | 8.7% | 5.3% |
| Avg. Similarity | 0.39 | 0.38 | 0.31 |
5.3.2 Source of Empirical Gains
Since LLAMA2 has been pre-trained on some evaluation benchmarks, it remains questionable whether the empirical gains in downstream tasks are due to knowledge overlap between distillations from LLAMA2 and the evaluation benchmarks. To this extent, we further demonstrate that CANDLE distilled models perform better due to improved generalizability rather than relying on data overlap with the evaluation data. We use SentenceBERT Reimers and Gurevych (2019) to measure the textual similarity between the distilled knowledge and the evaluation data for each task. We then calculate the ratio of data that exhibits semantic overlap with a similarity score exceeding 0.5 and also report the average similarity. The results are shown in Table 6. Based on the results, we observe that the distilled knowledge has minimal overlap with the evaluation set. This indicates that the empirical gain primarily stems from our distilled knowledge, which improves the generalizability of the models, rather than relying on knowledge overlap with the evaluation sets.
@@ 6 Conclusion
This paper introduces CANDLE, a distillation framework that realizes the chain of conceptualization and instantiation over CSKBs. We demonstrate the efficacy of CANDLE through comprehensive evaluations of the distilled knowledge and its positive impact on downstream tasks. Our research sheds light on distilling LLMs to enable more robust and generalizable commonsense reasoning.
@@
The major limitation of CANDLE lies in the significant cost of distilling LLMs to obtain substantial knowledge. While the instantiation step of CANDLE utilizes the open foundation model LLAMA2, the conceptualization is still performed by ChatGPT due to the unsatisfactory performance of other open-source LLMs and the high quality of ChatGPT’s generation. Consequently, a considerable amount of funding is required to distill conceptualizations for CANDLE to function effectively. Future works should address this issue by exploring advanced prompting methods or employing stronger open-source LLMs as the foundation for distilling conceptualizations.
Furthermore, it should be noted that CANDLE has only been validated on ATOMIC. However, CANDLE is not limited to any specific format of commonsense knowledge, allowing it to operate on any CSKB. Future research can address this by extending the evaluation of CANDLE to other CSKBs and conducting follow-up experiments to explore their benefits on more downstream tasks.
Another interesting direction to investigate is utilizing the chain of conceptualization and instantiation as a foundation for enhancing weak-to-strong generalization Burns et al. (2023). By conceptualizing and instantiating weak supervision data, we can generate more robust and generalized training signals, which ultimately strengthens the learning process. This can also be effectively incorporated into the training process of self-rewarding language models Yuan et al. (2024).
@@
To avoid generating harmful or unethical content from LLMs like ChatGPT and LLAMA2, we recruit four expert annotators, who are graduate and undergraduate students specializing in machine commonsense in natural language processing, to verify the ethics and potential harm of the generated content. A thorough assessment of a random sample has been conducted, and no significant harm has been identified. All training and evaluation datasets used are publicly available and shared under open-access licenses solely for research purposes, aligning with their intended usage. These datasets have been carefully anonymized and desensitized to protect data privacy and confidentiality. The expert annotators involved in this study are fully aware of the annotation protocol and the intended use of their annotations. Their participation in this research is voluntary, and they have agreed to contribute without receiving any compensation. Thus, the authors believe that this paper does not raise any ethical concerns.